If thrice of Aman's age 6 years ago be subtracted from twice his present age the result would be equal to his present age.
Find Aman's present age.

Found 2 solutions by stanbon, advanced_Learner:
Answer by stanbon(75887)

If thrice of Aman's age 6 years ago be subtracted from twice his present age the result would be equal to his present age.
Find Aman's present age.
-----
Equation:
.
2a - 3(a-6) = a
------
2a - 3a + 18 = a
----
2a = 18
a = 9 yrs (Aman's age now)
--------------
Cheers,
Stan H.
---------

Answer by advanced_Learner(501)

2x-3(x-6)=x
x=9
check
present age 9
twice present age 18
6 years ago 3
thrice 6 years 9
18-9=9